# DiffVLA++: Bridging Cognitive Reasoning and End-to-End Driving through Metric-Guided Alignment

### v3

@@ 1 Introduction
End-to-end (E2E) autonomous driving frameworks have made significant progress in recent years by directly mapping multi-modal sensory inputs to control outputs or trajectories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. These models benefit from strong spatio-temporal modeling capabilities and are trained to produce physically plausible driving behaviors under normal operating conditions. However, they often struggle in long-tail scenarios due to their limited capacity for high-level scene understanding and semantic reasoning [14, 15, 16, 17]. This weakness stems from their reliance on structured pattern recognition modules rather than human level cognitive modeling, leaving them devoid of generalizable world knowledge that human drivers naturally exploit.
To address this limitation, recent work has explored Large Language Models (LLMs) and Vision-Language-Action (VLA) models in autonomous driving. Some studies using LLMs [18, 19, 20] rely on their extensive world knowledge to generate high-level planning decisions, while others [21, 22, 23, 24, 25] directly generate trajectories and thus constitute Vision-Language-Action (VLA) models. In DiffVLA++, we aim to leverage the VLA model to generate semantically rich, directly usable driving trajectories.
@@ 3 Conventional End-to-End Driving Model
The conventional end-to-end (E2E) driving module in our framework operates on a dense Bird’s Eye View (BEV) representation generated by BevFormer [27]. We adopt VoVNet-99 [28] as the image backbone to enhance the visual representational capacity. The BEV feature map is discretized into a $128\times 128$ grid, covering a spatial extent of $64\times 64$ meters along the ego-vehicle’s $x$ and $y$ axes.

### v4

@@ 1 Introduction
End-to-end (E2E) autonomous driving frameworks have achieved remarkable progress in recent years by directly mapping multi-modal sensory inputs to control signals or driving trajectories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. These models benefit from powerful spatio-temporal representations and can generate physically plausible driving behaviors under normal conditions. However, they often struggle in long-tail or unseen scenarios due to their limited capability in high-level scene understanding and semantic reasoning [14, 15, 16, 17].
DiffVLA [18] attempts to enhance the reliability of spatio-temporal reasoning for both static and dynamic objects by integrating dense and sparse Bird’s Eye View (BEV) perception streams. Nevertheless, its limitation arises from the heavy reliance on structured pattern-recognition modules rather than human-level cognitive modeling, which leaves it lacking the generalizable world knowledge that human drivers naturally possess. In DiffVLA++, we retain only the dense BEV branch and aim to address this issue by incorporating cognitive knowledge to enhance reasoning ability.
To further tackle this limitation, recent studies have explored the integration of Large Language Models (LLMs) and Vision-Language-Action (VLA) architectures into autonomous driving systems. Some approaches [19, 20, 21] leverage the extensive world knowledge encoded in LLMs to generate high-level driving decisions, while others [22, 23, 24, 25, 26] directly produce driving trajectories, forming Vision-Language-Action (VLA) models. In DiffVLA, we adopt the former paradigm for its simplicity to integration. In DiffVLA++, we further exploit the VLA framework to generate semantically rich and directly executable driving trajectories.
@@ 3 Conventional End-to-End Driving Model
The conventional end-to-end (E2E) driving module in our framework operates on a dense BEV representation generated by BevFormer [28]. We adopt VoVNet-99 [29] as the image backbone to enhance the visual representational capacity. The BEV feature map is discretized into a $128\times 128$ grid, covering a spatial extent of $64\times 64$ meters along the ego-vehicle’s $x$ and $y$ axes.